OpenStudy (josee):

lim[x:1,(√(2-x)-1)/(√(5-x)-2)]

OpenStudy (josee):

ohhh i already did that one

OpenStudy (josee):

can you help me ith another one? im having trouble with cube root

OpenStudy (josee):

lim of x to 1 of \[(\sqrt[3]{x}-1)/[(\sqrt{x}-1)\]

OpenStudy (zarkon):

\[\frac{\sqrt[3]{x}-1}{\sqrt{x}-1}=\frac{\sqrt[3]{x}-1}{\sqrt{x}-1}\frac{\sqrt{x}+1}{\sqrt{x}+1}\frac{x^{2/3}+x^{1/3}+1}{x^{2/3}+x^{1/3}+1}\]

OpenStudy (zarkon):

use \[x^3-1=(x-1)(x^2+x+1)\]